OpenStudy (nuccioreggie):

Convert 0.0000001 to a power of 10. 107 10−7 106 10−6

OpenStudy (anonymous):

Well 2 of them are obviously not it.

OpenStudy (nuccioreggie):

Yes

OpenStudy (mathstudent55):

Count the number of decimal places in the number. Then make that number negative. That is the exponent on the 10.

OpenStudy (sachintha):

1/10 = 0.1 So for 0.0000001 = ?

OpenStudy (sachintha):

\[\frac{1}{10^2}=\frac{1}{100}=0.01\]

OpenStudy (nuccioreggie):

I got B

OpenStudy (mathstudent55):

Correct.

OpenStudy (sachintha):

Well Done!